OpenStudy (anonymous):

which is equivalent to spr 45 = 8 sqr 5 show work please

OpenStudy (anonymous):

which what? >.>

OpenStudy (anonymous):

spr = square root?

OpenStudy (anonymous):

\[Isn't\ a\ stupid\ question\ ?\]

OpenStudy (anonymous):

yes sqr = square root symbol

OpenStudy (smurfy14):

answer choices plz

OpenStudy (anonymous):

\[none \ of\ the\ above\]

OpenStudy (anonymous):

sqr 45 = sqr9* sqr 5 = 3* sqr 5

OpenStudy (anonymous):

\[\sqrt{45}= 3 \sqrt{5}\]

OpenStudy (anonymous):

answer choices r: 5 sqr 11 11 sqr 5 8 sqr 50 17 sqr 5